Tesla Corporation needs to raise funds to finance a plant expansion, and it has decided to issue 20-year zero coupon bonds to raise the money. The required return on the bonds will be 8 percent. Assume a par value of $1,000 and semiannual compounding.

a.

What will these bonds sell for at issuance? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  Issue price $   
b.

Using the IRS amortization rule, what interest deduction can the company take on these bonds in the first year? In the last year? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Interest deduction
  First year $   
  Last year $   
c.

Using the straight-line method, what interest deduction can the company take on these bonds in the first year? In the last year?. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Interest deduction
  First year $   
  Last year $   
d.

Based on your answers in (b) and (c), which interest deduction method would the company prefer?

IRS amortization rule
Straight-line method
Answer #1

a)

Price = FV / (1 + (r / 2))^t x compounding rate

Price = $1,000 / (1 + (8% / 2))^20 x 2

Price = $ 208.29

Issue price $ 208.29

b)

Bond’s value for the year.

Value beginning of 1st Year = $208.29

Value at end of 1st Year = FV / (1+ (r / 2))^Remaining Years x Compounding Rate

Value at end of 1st Year = $1,000 / (1 + (8% / 2))^19 x 2

*Remember that 1 year has past thus the 24 in the equation.

Value at end of 1st Year = $225.29

Hence, interest deduction in the first year is $225.29 - $208.29 = $17

Value at Beginning of Last Year = FV / (1 + (r / 2))^Remaining Time x Compound Time

Value at Beginning of Last Year = $1,000 / (1 + (8% / 2))^1 x 2

*Remember that 19 years have passed now and there is only 1 year left until maturity

Value at Beginning of Last Year = $924.56

Value at End of Last Year = $1,000 *Remember $1,000 is always the FV

Interest Deduction = Value of Bond at End of Year – Value of Bond at Beginning of Year

Interest Deduction = $1,000 - $924.56

Interest Deduction = $75.44

Interest deduction
  First year $ 17  
  Last year $ 75.44

c)

Previous IRS regulations required a straight-line calculation of interest.

The total interest received by the bond holder is:

Total interest = $1,000 – 208.29 = $791.71

The annual interest deduction is simply the total interest divided by the maturity of the bond, so the straight-line deduction is:Annual interest deduction = $791.71 / 20 = $39.59

d)

The company will prefer straight-line methods when allowed because the valuable interest deductions occur earlier in the life of the bond

Straight-line method